If the engine of a car of mass m is capable of power P, the minimum time needed for the car to accelerate from rest to a speed v is?


1
Expert's answer
2021-08-17T17:27:34-0400

Gives

Power =P

Speed=v

Only kinetic energy existing potential energy zero

∆E=K.E+P.E

P.E=0

∆E=K

"∆E=\frac{1}{2}mv^2"

Power

"P=\frac{∆E}{t}"

∆E=Pt

"\frac{1}{2}mv^2=Pt"

"t=\frac{mv^2}{2P}"